OpenStudy (anonymous):

find the derivative of 2x -5sqrt(x) using rules of differentiations please

OpenStudy (anonymous):

2 - 10x

OpenStudy (anonymous):

can you show steps pls

OpenStudy (anonymous):

yeah sure, \[\frac{ d}{ dx }(2x-5x ^{2})\] \[=\frac{ d}{ dx }(2x)-\frac{ d}{ dx }(5x ^{2})\] \[=2 - 5\frac{ d }{ dx }(x ^{2})\] \[=2 - 5 * 2x\] \[= 2- 5x\]

OpenStudy (anonymous):

ok thx very much. i have another one similar f(t)= 2t^2 - sqrt(t^3)

OpenStudy (anonymous):

its quite similar, you have to do same thing. simply remember, while differentiating some variable exponent some constant number, you just have to multiply the exponent to coefficient of the term and then write the variable after decreasing the exponent by one. Do this for each term in an expression and you would be done :) as for your ques, \[f'(x) = \frac{ d }{ dx }(2t ^{2} - \sqrt{t ^{3}})\] \[= \frac{ d }{ dx }(2t ^{2}) - \frac{ d }{ dx }(\sqrt{t ^{3}})\] \[=2 \frac{ d }{ dx }(t ^{2}) - \frac{ d }{ dx }(t ^{\frac{ 3 }{ 2 }})\] \[=2 *2t - \frac{ 3 }{ 2 }(t ^{\frac{ 1 }{ 2 }})\] \[=4t - \frac{ 3 }{ 2 }(\sqrt{t})\]

OpenStudy (anonymous):

wow thanks very much. i have few other questions, dont know if you would mind helping me?